a projectile is fired in such a way that its horizontal range is equal to three times its maximum hight. what is the angle of projection?
Answer #1
gt = 0
t = Vyi/g. The peak altitude is Vyit - .5gt^2 The total flight time is twice t since the time until we reach the peak altitude is equal to half the timefrom the peak altitude to ground. The horizontal range is Dx = Vix * 2t
3Vyit-1.5gt^2 = Vix*2t. If the angle of projection is a tan(a) = Vyi/Vx, so Vxi*tan(a) = Vyi
3Vyi-1.5gt=2*Vix. Substitute expression for t = Vyi/g
3Vyi-1.5gVyi/g=2*Vix, so Vyi[3-1.5] = 2Vix. Therefore Vyi/Vxi = tan(a) = 4/3 The angle of projection is then 53.13 degrees
answered by: Thabang